A population consists of three numbers (3,4,7). Consider all possible samples of size 2 which can be drawn from the population. Find the variance of the sampling distribution of the sample means.


1
Expert's answer
2022-03-28T15:55:23-0400

"\mu(3,4)=(3+4)/2=3.5"

"\mu(3,7)=(3+7)/2=5"

"\mu(4,7)=(4+7)/2=5.5"

f(3.5)=f(5)=f(5.5)=1/3

"E(x)=\sum fx=1/3(3.5+5+5.5)=4.7"

"\sigma^2=\sum fx^2-(\sum fx)^2=1/3(12.25+25+30.25)-22.09=0.41"